# Diffusion-based Dynamic Contract for Federated AI Agent Construction in Mobile Metaverses

### v1

@@ IV-D Optimal Two-Period Contract
The number of variables in the dynamic contract optimization problem (43) has been reduced from $4K$ to $2K$. Currently, researchers rely on heuristic algorithms to solve this problem [22, 16]. However, in practical scenarios, the parameters of ESs and the cloud server, such as communication channel conditions, fluctuate dynamically. As a result, heuristic algorithms must be frequently re-designed and re-implemented, leading to substantial computing resource consumption. Due to their adaptability to dynamic environments, DRL algorithms have been successfully applied for optimal static contract design [24, 27]. Notably, researchers have adopted GDMs to identify optimal static contracts, achieving superior performance compared with traditional DRL algorithms [19, 29]. Building on this foundation, we propose the EDMSAC algorithm that leverages GDMs to generate optimal dynamic contracts, effectively adapting to the changing environment during federated AI agent construction.
@@ V-A Markov Decision Process Formulation
MDP serves as a mathematical framework for modeling decision-making problems. It is represented as a five-element tuple consisting of a state space, an action space, a reward function, a discount factor determining the weight of the future reward relative to the current reward, and state transition probabilities. In the following, we explicitly define the state space, action space, and reward function of the MDP framework.
@@ V-A1 State space
where $N$ and $K$ are constants, while the other terms vary with each training round. Notably, the size of the state space $\bm{S}$ is $(6+3K+K^{2})$.
@@ V-A3 Reward function
where $\lambda$ is a pre-defined parameter that controls the level of the action penalty.
1
2Initialize policy network $\pi_{\bm{\theta}}$, critic networks $Q_{\bm{\upsilon}_{1}},Q_{\bm{\upsilon}_{2}}$, target networks $\hat{\pi}_{\hat{\bm{\theta}}},Q_{\hat{\bm{\upsilon}}_{1}},Q_{\hat{\bm{\upsilon}}_{2}}$, relay buffer $\mathcal{D}$, pruning rate $\varrho$, and mask matrix $\bm{W}$.
3## Training
4for the training step $v=1$ to $V$ do
5       for the number of transitions $z=1$ to $Z$ do
6             ### Generating contracts
9            Store record $(\bm{S}_{z},\bm{A}_{0},\bm{S}_{z+1},r_{z})$ into $\mathcal{D}$.
10            Sample a random mini-batch of transitions $\mathcal{B}_{z}$ with the size $B$ from $\mathcal{D}$.
11            ### Dynamic pruning
13            Use the mask matrix $\bm{W}$ to mask output neurons based on the pruning rate $\varrho$ by (58).
14            ### Parameter updates
15            Update the policy $\pi_{\bm{\theta}}$ using $\mathcal{B}_{z}$ by (59).
18       end for
19
20 end for
21Reconstruct the compact policy networks.
22## Inference
23Input a state $\bm{S}$.
return $\bm{A}_{0}=\Omega^{\star}$.
Algorithm 1 EDMSAC for Dynamic Contract Design
@@ V-C Enhanced Diffusion Policy for Dynamic Contract Design
where the end sample of the reverse chain is $\tilde{\bm{A}}_{0}$. According to [19, 34], $g_{\bm{\theta}}(\tilde{\bm{A}}_{m-1}|\tilde{\bm{A}}_{m},\bm{S})$ is modeled as a Gaussian distribution $\mathcal{N}(\tilde{\bm{A}}_{m-1};\mu_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m),\Sigma_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m))$ based on (49), with the covariance matrix $\Sigma_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)=\zeta_{m}\mathbf{I}$ and the mean $\mu_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)$ constructed as
$\mu_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)=\frac{1}{\sqrt{\chi_{m}}}\bigg{(}\tilde{\bm{A}}_{m}-\frac{\zeta_{m}}{\sqrt{1-\bar{\chi}_{m}}}\xi_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)\bigg{)},$ (53)
@@
Figure 4: Two case studies of mobile phone operations using an AI agent. We call the Qwen-VL-Plus and Qwen-VL-Max models through the Qwen-VL API to construct the mobile AI agent, utilizing Android Debug Bridge (ADB) to operate a mobile phone with the Android OS system, where ADB can simulate all operations of the AI agent [30].
(a) Case 1.
(b) Case 2.
Figure 5: Performance evaluation of mobile AI agents.
To enhance the denoising efficiency, we employ dynamic structured pruning techniques to linear layers of $\xi_{\bm{\theta}}(\cdot,\cdot,\cdot)$. Considering $\xi_{\bm{\theta}}(\cdot,\cdot,\cdot)$ with $H$ fully connected layers, we denote the parameters in the $h$-th fully connected layer as $\bm{\theta}^{(h)}\in\mathbb{R}^{x\times y},\>h\in\{1,\ldots,H\}$. We first calculate the importance of each row in $\bm{\theta}^{(h)}$, which is expressed as
$||\bm{\theta}^{(h)}_{i}||_{2}=\sqrt{\sum_{j=1}^{y}\theta_{i,j}^{2}},\>i=1,2,\ldots,x.$ (57)
Then, we introduce a mask matrix $\bm{W}^{(h)}\in\mathbb{R}^{x\times y}$ to prune fully connected layers with low importance based on the pre-defined pruning rate. The elements of $\bm{W}^{(h)}\in\mathbb{R}^{x\times y}$ are binary, where $\bm{W}^{(h)}_{x}=0$ indicates that the output neuron $o_{x}^{(h)}$ is pruned, while $o_{x}^{(h)}$ remains otherwise. Hence, the pruned parameters in the $h$-th fully connected layer are given by
@@
Algorithm 1 illustrates the procedure of the EDMSAC algorithm, where its computational complexity primarily consists of three components: In action sampling, the computational overhead arises from environment interaction and the reverse process, with a complexity of $\mathcal{O}(VZ(1+M|\bm{\theta}|))$ [36]. In dynamic pruning, the computational overhead is due to the calculations of neuron importance and masking, leading to a complexity of $\mathcal{O}(\sum_{h=1}^{H}(x^{(h)}y^{(h)}+x^{(h)}\log x^{(h)}))$ [37]. In parameter updates, the computational complexity stems from updating network parameters, given by $\mathcal{O}(V(B+1)(|\bm{\theta}|+|\bm{\upsilon}|))$ [36]. Therefore, the computational complexity of the EDMSAC algorithm is approximately $\mathcal{O}(VZ(1+M|\bm{\theta}|)+\sum_{h=1}^{H}(x^{(h)}y^{(h)}+x^{(h)}\log x^{(h)})+V(B+1)(|\bm{\theta}|+|\bm{\upsilon}|))$.
TABLE I: Simulation Parameters
@@
Number of ESs ($N$) $\{3,6,12,18\}$
Pre-defined hyperparameters ($\delta,L,\varepsilon,\lambda$) $\{0.02,8,2,0.01\}$
Constant related to the hardware architecture of ES $n$ ($\varsigma_{n}$) $10^{-23}$
Number of CPU cycles required for perceiving feature data ($\Upsilon_{n}^{Per}$) $100\>\rm{cycles/bit}$
CPU speed of ES $n$ ($f_{n}$) $64\>\rm{MHz}$
Size of feature data perceived by ES $n$ ($D_{n}^{Fea}$) $1\>\rm{MB}$
Effective switched capacitance of ES $n$ ($\rho_{n}$) $10^{-16}$
Number of CPU cycles required for creating one bit of agent modules ($\Upsilon_{n}^{Cre}$) $120\>\rm{cycles/bit}$
Transmit power of ES $n$ ($P_{n}$) $[20,33]\>\rm{dBm}$
Model size of agent modules constructed by ES $n$ ($D_{n}^{Agent}$) $10\>\rm{MB}$
Transmission rate from ES $n$ to the cloud server through the fiber link ($r_{n,c}$) $[1,3]\>\rm{Mbps}$
Unit cost of energy consumption ($\sigma$) $[0.5,1]$
Scaling factor that affects the overall magnitude of economic benefits ($\alpha$) $\{200,250\}$
Energy consumption of global AI agent integration ($E_{C}$) $[20,25]\>\rm{Joules}$
Discount factor ($\beta$) $\{0,0.5,1\}$
(a) EDMSAC vs PPO, SAC, and DDPG
(b) EDMSAC vs DMSAC and DMDDPG
(c) Performance results

### v2

@@ IV-D Optimal Two-Period Contract
The number of variables in the dynamic contract optimization problem (27) has been reduced from $4K$ to $2K$. Currently, researchers rely on traditional optimization algorithms, such as heuristic algorithms, to solve this problem [7, 14]. However, in practical scenarios, the parameters of ESs and the cloud server, such as communication channel conditions, fluctuate dynamically. As a result, traditional optimization algorithms must be frequently redesigned and re-implemented, leading to substantial computing resource consumption. Furthermore, traditional algorithms often fail to generalize to practical scenarios when the environment lacks complete and accurate information [29]. DRL algorithms have been successfully applied for optimal static contract design due to their adaptability to dynamic environments [41, 16]. In particular, PPO, as an on-policy and model-free algorithm, iteratively updates the policy towards optimal performance of contract generation by using a clipped surrogate objective [32], and SAC, as an off-policy algorithm, trains a stochastic policy for optimal dynamic contract generation by maximizing both the expected cumulative reward and policy entropy [31]. Notably, researchers have adopted GDMs to identify optimal static contracts, achieving superior performance compared with traditional DRL algorithms [31, 29]. In particular, the DMSAC algorithm can improve stability and performance over baseline DRL algorithms [4]. Building on this foundation [4], we propose the EDMSAC algorithm that leverages lightweight GDMs to generate optimal dynamic contracts, effectively adapting to the changing environment during federated AI agent construction.
@@ V-A Markov Decision Process Formulation
MDP is a mathematical framework for modeling decision-making problems [17], which is represented as a five-element tuple consisting of a state space, an action space, a reward function, a discount factor determining the weight of the future reward relative to the current reward, and state transition probabilities. In the following, we explicitly define the state space, action space, and reward function of the MDP framework.
@@ V-A1 State space
where $N$ and $K$ are constants that do not evolve over time, while the other terms vary with each training round. Notably, the size of the state space $\bm{S}$ is $(6+3K+K^{2})$.
@@ V-A3 Reward function
where $\lambda$ is a pre-defined parameter that controls the degree of the action penalty.
@@ V-B Denoising Diffusion Probabilistic Models
1
2Initialize policy network $\pi_{\bm{\theta}}$, critic networks $Q_{\bm{\upsilon}_{1}},Q_{\bm{\upsilon}_{2}}$, target networks $\hat{\pi}_{\hat{\bm{\theta}}},Q_{\hat{\bm{\upsilon}}_{1}},Q_{\hat{\bm{\upsilon}}_{2}}$, relay buffer $\mathcal{D}$, pruning rate $\varrho$, and mask matrices $\bm{W}$.
3## Training
4for the training step $v=1$ to $V$ do
5 for the number of transitions $z=1$ to $Z$ do
6    ### Generating contracts
9     Store record $(\bm{S}_{z},\bm{A}_{0},\bm{S}_{z+1},r_{z})$ into $\mathcal{D}$.
10   end for
11
12 ### Dynamic pruning
14  Use the mask matrices $\bm{W}$ to mask output neurons based on the pruning rate $\varrho$ by (42).
15 ### Parameter updates
16  Sample a random mini-batch of transitions $\mathcal{B}_{z}$ with size $B$ from $\mathcal{D}$.
17  Update the policy $\pi_{\bm{\theta}}$ using $\mathcal{B}_{z}$ by (43).
20 end for
21Reconstruct the compact policy networks.
22## Inference
23Input a state $\bm{S}$.
return $\bm{A}_{0}=\Omega^{\star}$.
Algorithm 1 EDMSAC for Dynamic Contract Design
@@ V-C Enhanced Diffusion Policy for Dynamic Contract Design
where the end sample of the reverse chain is $\tilde{\bm{A}}_{0}$. According to [28], $g_{\bm{\theta}}(\tilde{\bm{A}}_{m-1}|\tilde{\bm{A}}_{m},\bm{S})$ is modeled as a Gaussian distribution $\mathcal{N}(\tilde{\bm{A}}_{m-1};\mu_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m),\Sigma_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m))$ based on (33), with the covariance matrix $\Sigma_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)=\frac{\zeta_{m}(1-\bar{\chi}_{m-1})}{1-\bar{\chi}_{m}}\mathbf{I}$ and the mean $\mu_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)$ constructed as
$\mu_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)=\frac{1}{\sqrt{\chi_{m}}}\bigg(\tilde{\bm{A}}_{m}-\frac{\zeta_{m}\tanh{\xi_{\bm{\theta}}(\tilde{\bm{A}}_{m},\bm{S},m)}}{\sqrt{1-\bar{\chi}_{m}}}\bigg),$ (37)
@@
To enhance denoising efficiency, we employ dynamic structured pruning techniques to linear layers of $\xi_{\bm{\theta}}(\cdot,\cdot,\cdot)$. Considering $\xi_{\bm{\theta}}(\cdot,\cdot,\cdot)$ with $H$ fully connected layers [29], we denote the parameters in the $h$-th fully connected layer as $\bm{\theta}^{(h)}\in\mathbb{R}^{x\times y},\>h\in\{1,\ldots,H\}$. We first calculate the importance of each row in $\bm{\theta}^{(h)}$, which is expressed as
$\bm{\theta}^{(h)}_{i}||_{2}=\sqrt{\sum_{j=1}^{y}\theta_{i,j}^{2}},\>i=1,2,\ldots,x.$ (41)
Then, we introduce mask matrices $\bm{W}^{(h)}\in\mathbb{R}^{x\times y}$ to prune fully connected layers with low importance based on the pre-defined pruning rate. The elements of $\bm{W}^{(h)}\in\mathbb{R}^{x\times y}$ are binary, where $\bm{W}^{(h)}_{x}=0$ indicates that the output neuron $o_{x}^{(h)}$ is pruned, while $o_{x}^{(h)}$ remains otherwise. Hence, the pruned parameters in the $h$-th fully connected layer are given by
@@
Figure 4: Two case studies of mobile phone operations based on an AI agent. We call the Qwen-VL-Plus and Qwen-VL-Max models through the Qwen-VL API to construct the mobile AI agent, utilizing Android Debug Bridge (ADB) to operate a mobile phone with the Android OS system, where ADB can simulate all operations of the AI agent [26].
(a) Case 1.
(b) Case 2.
Figure 5: Performance evaluation of mobile AI agents.
Algorithm 1 illustrates the procedure of the EDMSAC algorithm, where its computational complexity primarily consists of three components: In action sampling, the computational overhead arises from environment interaction and the reverse process, with a complexity of $\mathcal{O}(VZ(1+M|\bm{\theta}|))$ [4]. In dynamic pruning, the computational overhead is due to the calculations of neuron importance and masking, leading to a complexity of $\mathcal{O}(\sum_{h=1}^{H}(x^{(h)}y^{(h)}+x^{(h)}\log x^{(h)}))$ [11]. In parameter updates, the computational complexity stems from updating network parameters, given by $\mathcal{O}(V(B+1)(|\bm{\theta}|+|\bm{\upsilon}|))$ [4]. Therefore, the computational complexity of the EDMSAC algorithm is $\mathcal{O}(VZ(1+M|\bm{\theta}|)+\sum_{h=1}^{H}(x^{(h)}y^{(h)}+x^{(h)}\log x^{(h)})+V(B+1)(|\bm{\theta}|+|\bm{\upsilon}|))$.
VI Simulation Results
VI-A Experimental Setup
Experiment Settings: We consider ESs with homogeneous hardware capabilities and classify them into two types: high WTP and low WTP [31, 29]. Specifically, $\theta^{t}_{1}$ and $\theta^{t}_{2}$ are randomly sampled within $[15,20)$ and $[20,25)$, respectively [30]. $p^{1}_{k}$ and $p^{2}_{k}$ are randomly generated following the Dirichlet distribution [30]. The setting of simulation parameters is summarized in Table II. For the hyperparameter setting of the EDMSAC algorithm, DM-based actor networks adopt the variance proportional noise schedule strategy [4], where the noise schedule strategy determines the amount of noise injected into input data at each timestep during the forward diffusion process [29]. Both the actor and critic networks are trained using the Adam optimizer, with a weight decay rate of $0.0001$ to regularize model weights [4]. The learning rates are set to $2\times 10^{-7}$ for the actor networks and $2\times 10^{-6}$ for the critic networks, respectively [4]. In addition, the discount factor $\gamma$ is set to $0.99$ [5], the update rate of the target networks $\tau$ is set to $0.005$ [5], the temperature coefficient $\kappa$ is set to $0.05$ [5], the denoising step $M$ is set to $6$ [4], and the pruning rate $\varrho$ is set to $0.5$ [29]. Our experiments use PyTorch with CUDA 12.0 on an NVIDIA GeForce RTX 3080 Laptop GPU.
TABLE II: Simulation Parameters
@@ VI-A Experimental Setup
Number of ESs (NN[35] $\{3,6,12,18\}$
Pre-defined hyperparameters ($\delta,L,\varepsilon,\lambda$[1] $\{0.02,8,2,0.01\}$
Constant related to the hardware architecture of ES $n$ ($\varsigma_{n}$[3] $10^{-23}$
Number of CPU cycles required for perceiving feature data ($\Upsilon_{n}^{Per}$[1] $100\>\rm{cycles/bit}$
CPU speed of ES $n$ ($f_{n}$[1] $64\>\rm{MHz}$
Size of feature data perceived by ES $n$ ($D_{n}^{Fea}$[41] $1\>\rm{MB}$
Effective switched capacitance of ES $n$ ($\rho_{n}$[1] $10^{-16}$
Number of CPU cycles required for creating one bit of agent modules ($\Upsilon_{n}^{Cre}$[1] $120\>\rm{cycles/bit}$
Transmit power of ES $n$ ($P_{n}$[35] $20,33]\>\rm{dBm}$
Model size of agent modules constructed by ES $n$ ($D_{n}^{Agent}$[41] $10\>\rm{MB}$
Transmission rate from ES $n$ to the cloud server through the fiber link ($r_{n,c}$[1] $1,3]\>\rm{Mbps}$
Unit cost of energy consumption ($\sigma$[14] $[0.5,1]$
Scaling factor that affects the overall magnitude of economic benefits ($\alpha$[14] $\{200,250\}$
Energy consumption of global AI agent integration ($E_{C}$[1] $20,25]\>\rm{Joules}$
Discount factor ($\beta$[7] $\{0,0.5,1\}$
Algorithm Comparisons: To evaluate the performance of the proposed EDMSAC algorithm, we compare it with mainstream DRL algorithms in optimal contract design, and present the core hyperparameter settings used for each algorithm:
  • DMSAC [4]: The DMSAC algorithm does not incorporate the pruning rate. All hyperparameters are kept consistent with those used in the EDMSAC algorithm. It is worth noting that the computational complexity of the DMSAC algorithm is $\mathcal{O}(VZ(1+M|\bm{\theta}|)+V(B+1)(|\bm{\theta}|+|\bm{\upsilon}|))$.
  • DMDDPG [5]: We set the exploration noise to a fixed Gaussian distribution with zero mean and standard deviation of $0.1$. All other basic hyperparameters are the same as those used in the EDMSAC algorithm.
  • PPO [32]: We apply the $\tanh$ activation function as the action bounding method, and set the discount factor $\gamma$ and the learning rate to $0.95$ and $3\times 10^{-5}$, respectively. All other basic hyperparameters are default, which are commonly used in prior literature [32].
  • SAC [31]: We set the learning rate for the temperature coefficient $\kappa$ to $10^{-4}$. The learning rates are set to $10^{-4}$ for the actor networks and $10^{-3}$ for the critic networks, respectively. All other basic hyperparameters are the same as those used in the EDMSAC algorithm.
  • DDPG [29]: We set the learning rates to $10^{-4}$ for the actor networks and $10^{-3}$ for the critic networks, respectively. The discount factor $\gamma$ is set to $0.99$. All other basic hyperparameters are the same as those used in the DMDDPG algorithm.
Scheme Comparisons: To evaluate the performance of the proposed dynamic contract scheme, we compare it with other typical contract schemes:
(a) EDMSAC vs PPO, SAC, and DDPG.
(b) EDMSAC vs DMSAC and DMDDPG.
(c) Performance results.
@@
Figure 7: The resource consumption of the proposed EDMSAC algorithm for performing one training and multiple inferences.
VI-B Practical Implementation Guidelines
We present a practical implementation guideline for the proposed federated AI agent construction framework, investigating how the proposed framework can be closely integrated with the proposed dynamic contract scheme and the EDMSAC algorithm in practical scenarios.
  • Infrastructure Initialization: Each ES is equipped with sufficient computing resources (e.g., embedded GPUs) and preloaded with basic AI models from the cloud server. Upon boot-up, ESs establish low-latency communication with the cloud server through fiber links.
  • Dynamic Contract Generation: Based on the predicted WTP levels of ESs, the cloud server executes the proposed EDMSAC algorithm to generate the optimal two-period dynamic contract. Then, the cloud server releases the generated dynamic contract to ESs at the beginning of federated AI agent construction.
  • Agent Module Construction: The ESs first select their preferred contract item according to their WTP levels. Following contract confirmation, the ESs construct agent modules using perceived feature data. Then, the ESs upload constructed agent modules to the cloud server. The process can be referred to Section III-A.
  • User Interaction Interfaces: The cloud server integrates uploaded modules into complete AI agents and deploys them to ESs. Users can interact with them through intuitive mobile-based interfaces, such as a customized application connected via local IP networks, enabling real-time AI agent services, including personal assistants.
VI-C Robustness and Security Analysis
Due to its contract-theoretic design and EDMSAC-based optimization, the proposed scheme provides intrinsic robustness and security properties:
  1. 1)
    The IC and IIC constraints of the proposed scheme inherently mitigate strategic misreporting and free-riding behaviors. Since ESs maximize their long-term utilities only by truthfully revealing their WTP levels, any deviation or misreporting leads to a strictly lower expected utility.
  2. 2)
    The contract realization process across two periods enhances robustness against opportunistic behaviors. The cloud server issues contract-based rewards to the ESs only after successful agent module integration in each period, which discourages non-cooperative actions such as incomplete training.
  3. 3)
    Uncertainties arising from environments are reflected in the environmental state. The EDMSAC algorithm learns a contract policy that maximizes expected long-term utility, which exhibits robustness to moderate environmental perturbations without requiring explicit failure modeling.
VI-D Performance Evaluation of Mobile AI Agents
As illustrated in Fig. 4, we present two case studies in which a mobile AI agent operates a real smartphone to perform tasks. This experiment involves the distributed execution of agent modules, as well as task decomposition and coordination during execution. For the perception module, we leverage Qwen-VL-Plus to create it, and we use the ConvNextViT-document model and the GroundingDINO model as text and icon recognition tools, respectively [26]. For the decision-making, reflection, and planning modules, we leverage Qwen-VL-Max to create them. To evaluate the performance of the constructed AI agent, we adopt several evaluation metrics: Decision Accuracy (DA), which measures the accuracy of the decision-making module; Reflection Accuracy (RA), which assesses the accuracy of the reflection module; and Planning Accuracy (PA), which evaluates the accuracy of the planning module. As shown in Fig. 5, the performance of the AI agent in executing simple tasks, such as Case 1, is superior to that in handling more complex tasks, as in Case 2, as evidenced by higher values of DA, RA, and PA. Notably, when executing simple tasks, the AI agent may perform additional rounds to verify task completion, leading to increased service latency.
VI-E Performance Evaluation of the EDMSAC Algorithm
In Fig. 6, we present a performance comparison between the proposed EDMSAC algorithm and several representative DRL algorithms in optimal contract design, including DMSAC, DMDDPG, PPO, SAC, and DDPG. From Figs. 6(a) and 6(b), we observe that the proposed EDMSAC algorithm outperforms these DRL algorithms in optimal dynamic contract design because DRL algorithms struggle in different suboptimal solution spaces, especially PPO algorithms. The superior performance of the proposed EDMSAC algorithm can be primarily attributed to two key reasons. First, the EDMSAC algorithm has a strong capability to capture complex environmental information because of GDMs [5], which helps avoid the identification of suboptimal dynamic contracts. Moreover, its fine-grained policy tuning mitigates the impact of environmental randomness, enhancing decision robustness. Second, through structural pruning techniques, the unimportant neurons in the DM-based actor networks are masked rather than permanently removed, thereby preserving the network structure while reducing computational redundancy. Therefore, the EDMSAC algorithm can achieve higher denoising efficiency, enabling the generation of more optimal and stable dynamic contracts. From Fig. 6(c), we also observe that the EDMSAC algorithm achieves a higher total profit of the cloud server compared with other DRL algorithms. Overall, the above analysis demonstrates the effectiveness and superiority of the proposed EDMSAC algorithm in optimizing dynamic contract design, which enhances the system performance of federated AI agent construction under dynamic information asymmetry.
In Fig. 7, we employ the Python package named CodeCarbon [^2] https://github.com/mlco2/codecarbon to estimate the resource consumption of the proposed EDMSAC algorithm for one training session and $20$ inference runs. The evaluation considers five dimensions: time ($\rm{h}$), $\text{CO}_{\text{2}}$ ($\rm{kg}$), CPU ($\rm{kWh}$), GPU ($\rm{kWh}$), and random access memory ($\rm{kWh}$). To enhance clarity, we scale the inference results of all five metrics by a factor of $1000$. We observe that all metrics for one training session are significantly higher than those for $20$ inference sessions. Especially, the total carbon emissions for performing $20$ inferences are only $0.123\>\rm{g}$. The reason is that the structural pruning technique can improve computational efficiency of the denoising process, thereby reducing total carbon emissions. These findings indicate the environmental sustainability of the proposed EDMSAC algorithm in optimal dynamic contract design.